OpenStudy (anonymous):

whats the answer to 0.2(7+2t)=0.4t+1.4

OpenStudy (anonymous):

7/5+.4t=.4t+1.4

OpenStudy (anonymous):

1.4-1.4=.4t-.4t

OpenStudy (anonymous):

no answer

OpenStudy (anonymous):

or no solution

OpenStudy (anonymous):

ya just no solution

OpenStudy (anonymous):

okkhow bout 0.5x-1.7=0,3x-1

OpenStudy (anonymous):

or no solution

OpenStudy (anonymous):

no answer

OpenStudy (anonymous):

Q: what is the answer to 0.2(7+2t)=0.4t+1.4 That answer is -17.5 and here is why 0.2(7+2t)=0.4t+1.4 7+2t = 0.4t+1.4/0.2 now 2t-0.4t = 1.4-7/0.2 now 1.6t = -5.6/0.2 now 1.6t = -28 now t = -28/1.6 now t = -17.5

OpenStudy (anonymous):

its basic algebra

OpenStudy (anonymous):

To AICRAG: If it's basic algebra, you need to do the check at the end. 1. Try to plug your answer into original wquation and se if it's will come with a true statement. 2. On the second line, when you are dividing by 0.2, you need to divide both, not just the second one.